| Date | Comment |
| 20030808 - 20030809 |
Filter offsets wrong in filter database. On
20030808 focusing was done using the i filter. On 20030809 focusing was done
with the r filter. On 20030810 focus was done in every filter and a different
focus used for each filter. On 20030811 the filter database was corrected. |
| 20031004 |
Mistake found in code that calculated offset field positions. Fields with
-1 < declination < 0 are effected. Little data taken in this coordinate
range so far and no data that was uneffected by clouds. |
| 20031014 |
Coordinates for Wolf 1346 found to be wrong - corrected in the database
and object catalogue from 20031015 onward. |
| 20031111 - 20031112 |
Primary mirror removed from telescope and washed. Mirror support system
cleaned and checked. WFC CCD window cleaned and WFC pumped. Change in WFC
temperature servo electronics. |
| 20031113 |
Observations done with untested SDSU code that contains PCI extensions.
No effects on observations expected. |
| 20031113 - 20031114 |
Observations taken are missing the TCS header information. |
| 20031113 - ? |
Telescope oscillations sometimes effect the observing. |
| 20031113 - now |
WFC now runs with new temperature control code. |
| 20040531 - now |
Intermittent problems with the WFC focus mechanism. Some of these are
caused by network communication problems between the ICS and the TCS. This will
lead to the focus offset not beeing applied and one of the images being out of
focus. |
| 20050526 - now |
After aluminizing the image quality of the INT is not as good as it was
before. This is caused by the primary mirror not being in the correct position
relative to the field corrector. |